OpenStudy (anonymous):

Solve the following equation, explaining what you did in each step of your solution. After solving the equation, show the steps to check your solution. 3 x + 2(4 + 6 x ) = 113

OpenStudy (anonymous):

I found the answer but when I plugged it in, it wasn't right.

OpenStudy (jojokiw3):

Hm, expand 2(4+6x). What do you get?

OpenStudy (anonymous):

I get 2(4) + 2(6x)

OpenStudy (jojokiw3):

Which then equals?

OpenStudy (anonymous):

8 + 12x?

OpenStudy (jojokiw3):

Okay now replace 2(4+6x) with that an show me the entire equation.

OpenStudy (anonymous):

3x + 8 + 12x.

OpenStudy (jojokiw3):

= 113 right? So you now have.

OpenStudy (jojokiw3):

3x + 8 + 12x = 113. What do you do next?

OpenStudy (anonymous):

15x + 8 = 113?

OpenStudy (jojokiw3):

Okay, then what?

OpenStudy (anonymous):

We have to subtract 8 on both sides.

OpenStudy (jojokiw3):

Mhm, so what's the equation now??

OpenStudy (anonymous):

15x = 105.

OpenStudy (jojokiw3):

Mhm, so what does x =?

OpenStudy (anonymous):

7. But the thing is I plugged it in to the original equation and I ended up with a random number that's bigger than 113.

OpenStudy (jojokiw3):

Hm. Let's see. 3(7) + 2[4+6(7)] = 113. 21 + 2(4+42) = 113. 21 + 2(46) = 113. 21 + 92 = 113. 113 = 113.

OpenStudy (jojokiw3):

Seems alright to me. C:

OpenStudy (anonymous):

If you don't know what I did then this is what I did. 3(7) + 2(4) + 2(6*7) = 113 21 + 8 + 84 = 113 which is not 113.

OpenStudy (jojokiw3):

Lol 21 + 8 +84 does equal 113.

OpenStudy (jojokiw3):

8+84 = 92. 92 + 21 = 113.

OpenStudy (anonymous):

Okay I'm confused now.

OpenStudy (jojokiw3):

Lol 7 is the answer haha.

OpenStudy (anonymous):

Ugh! Thank You anyway. But why does math have to be so CONFUSING! Sorry but this is why I don't like math.

OpenStudy (jojokiw3):

Np.

OpenStudy (anonymous):

So thank you millions of times for undoing the confusion.

OpenStudy (anonymous):

Do you also do English?

OpenStudy (jojokiw3):

Yep.

OpenStudy (anonymous):

Oh. Great! Can you help me in that too? But since I'm kind of good a t math, I'll need help but not as much as math.

OpenStudy (jojokiw3):

Just tag me in the question.